### v1

@@ 3.2 Sensitivity
Our equilibrium concepts are designed to accommodate a variety of prior mechanisms developed for different settings and objectives. As a result, they do not require truth-telling to be a strict equilibrium. However, this leads to an apparent flaw: a trivial mechanism that always assigns Alice a score of 1 is technically truthful — and even SD-truthful — by definition. Clearly, such a mechanism fails to capture the essence of what a good scoring mechanism should achieve.
This highlights the need for an additional dimension in peer prediction mechanism design called the sensitivity (Zhang and Schoenebeck, 2023a). Intuitively, an effective scoring mechanism should meaningfully evaluate the information contained in agents’ reports and ensure that more informative reports receive higher rewards than uninformative ones.
To motivate sensitivity, we introduce effort into our model. On each question, Alice exerts effort with a probability of $e\in[0,1]$. If Alice exerts effort on a question, she will observe $X_{a}$ and report truthfully; otherwise, she will randomly sample a signal from the prior $\Pr(X_{a})$ and report it. Suppose Bob exerts effort on all questions and both Alice and Bob are truth-telling. Then, Alice’s score under mechanism $\mathcal{M}$ becomes a function of $e$, denoted as $S^{\mathcal{M}}(e)$.
Definition 3.
@@
Sensitivity measures how well a mechanism reflects changes in effort, and thus changes in information, in its expected score. The mechanism that consistently assigns 1 to all agents, regardless of their reports, has zero sensitivity since its expected score remains unchanged irrespective of effort (assuming $/0=0$ 0).
Sensitivity is closely related to other practical metrics of peer prediction mechanisms. Prior work shows that when scores are normally distributed, sensitivity is a sufficient statistic of budgetary efficiency: mechanisms with higher sensitivity can elicit a desired level of effort at lower budgetary costs (Zhang and Schoenebeck, 2023a). Furthermore, sensitivity is closely aligned with measurement integrity (Burrell and Schoenebeck, 2023), a concept used to assess peer prediction mechanisms’ ex-post fairness by measuring the correlation between the quality of agents’ reports and their scores allocated by the mechanism (Xu et al., 2024).
@@ 4 Prior Mechanisms
To the best of our knowledge, no existing peer prediction mechanism naturally achieves SD-truthfulness without imposing strong assumptions on the information structure. In this section, we introduce three simple and exemplary multi-task peer prediction mechanisms and explain why they are not naturally SD-truthful to motivate our methods.
The final score $S^{\mathcal{M}}$ of these mechanisms is usually the average of $K$ (sometimes independent) scores $S^{\mathcal{M}}_{i}$. We distinguish these two terms by calling $S^{\mathcal{M}}$ the final score and $S_{i}^{\mathcal{M}}$ the individual score.
@@ 4.1 Output Agreement (OA)
Given two agents’ reports on $n$ questions, a straightforward idea is to score Alice based on the fraction of questions on which both agents agree. This mechanism is commonly termed as output agreement (OA). In particular, OA repeatedly samples one of the $n$ questions (without replacement) and assigns Alice an individual score of 1 if both agents’ reports agree and 0 otherwise. The final score is the average of $n$ independent binary individual scores, i.e. $S^{OA}=\frac{1}{n}\sum_{i\in[n]}S_{i}^{OA}$ where $S_{i}^{OA}\in\{0,1\}$.
However, OA encourages agents to over-report the majority signal, which is truthful only under a special type of information structure.
Definition 4.
@@
The output agreement mechanism is truthful if agents’ signals are self-dominating.
Intuitively, self-dominating information structure implies that whenever Alice observes signal $\sigma$, she believes that Bob is most likely to also observe $\sigma$. Therefore, any manipulation will decrease the probability of agreeing with Bob. Furthermore, self-dominance is almost necessary for OA to be truthful: ignoring the case where two signals have the identical conditioned probability, if there exists $\sigma,\sigma^{\prime}$ such that $\Pr(X_{b}=\sigma|X_{a}=\sigma)<\Pr(X_{b}=\sigma^{\prime}|X_{a}=\sigma)$, then Alice is motivated to report $\sigma^{\prime}$ whenever she observes $\sigma$.
In fact, OA is also SD-truthful when signals are self-dominating. This is because each individual score $S^{OA}_{i}$ takes only two possible values, meaning that $S^{OA}_{i}(\tau)$ FOSD $S^{OA}_{i}(\theta)$ if and only if $\Pr(S^{OA}_{i}(\tau)=1)\geq\Pr(S^{OA}_{i}(\theta)=1)$. Furthermore, if multiple i.i.d. individual scores each satisfies SD-truthfulness, their average also satisfies SD-truthfulness (we will formally state this in Section 5.2). Consequently, for OA, SD-truthfulness reduces to truthfulness.
@@ 4.2 Peer Truth Serum (PTS)
Self-dominance is usually considered a strong assumption, especially when the signal space is large. (Radanovic et al., 2016) propose the Peer Truth Serum (PTS) mechanism which incentivizes truth-telling under more general information structures.
PTS assumes a symmetric and publicly known common prior $R(\sigma)=\Pr(X_{a}=\sigma)=\Pr(X_{b}=\sigma)$. The key idea is to weight the output agreement score reversely to the prior of the signal. The mechanism repeatedly samples a question $i$ without replacement and if there is an agreement $\hat{X}_{a,i}=\hat{X}_{b,i}=\sigma$, the mechanism assigns an individual score of 1 to Alice $S^{PTS}_{i}=1/R(\sigma)$, and assigns 0 if there is no agreement. The final score of PTS is the average of all $S^{PTS}_{i}$.
By re-weighting the agreement score, PTS relaxes the assumption of self-dominance to self-prediction. When the signal space is binary, self-prediction is equivalent to positive correlation, which is a strictly weaker assumption than self-dominance. However, self-prediction is not directly comparable with self-dominance when signals are non-binary.
Definition 5.
@@
The peer truth serum mechanism is truthful if agents’ signals are self-predicting.
Is PTS SD-truthful?  The short answer is no. Consider an individual score $S^{PTS}_{i}$, which can take at most $|\Sigma|+1$ distinct values. Let $\sigma_{min}$ be the least likely signal under the prior; then the largest score is $S_{max}=\max\left(S^{PTS}_{i}\right)=\frac{1}{R(\sigma_{min})}$. A necessary condition for PTS to be stochastic dominance-truthful is that truth-telling must maximize $\Pr\left(S^{PTS}_{i}(\tau)=S_{max}\right)$, which is the probability of Alice agreeing with Bob on signal $\sigma_{min}$.
However, this is not true. The probability of obtaining an individual score of $S_{max}$ under truth-telling is defined by the joint distribution $\Pr(X_{a}=\sigma_{min},X_{b}=\sigma_{min})$. If Alice plays an uninformative strategy $\mu$ which always reports $\sigma_{min}$, the probability of getting $S_{max}$ increases to $\Pr(X_{b}=\sigma_{min})$. Therefore, although $\mu$ decreases the expected score (because the mechanism is truthful), it leads to a score distribution that is not dominated by truth-telling, meaning that there exists an increasing utility function under which Alice strictly prefers $\mu$ to $\tau$.
@@ 4.3 Correlated Agreement (CA)
SD-truthfulness requires reasoning about the score distribution. Intuitively, mechanisms with a small number of possible scores are more likely to achieve SD-truthfulness. We now introduce the correlated agreement (CA) mechanism (Shnayder et al., 2016) whose score has three possible values: -1, 0, and 1. We show that this simple mechanism is not SD-truthful (for a different reason).
Suppose the joint distribution $\Pr(X_{a},X_{b})$ is known. The CA mechanism first computes the delta matrix $\Delta_{\sigma,\sigma^{\prime}}=\Pr(X_{a}=\sigma,X_{b}=\sigma^{\prime})-\Pr(X_{a}=\sigma)\Pr(X_{b}=\sigma^{\prime})$, which is the difference between the joint distribution and the product of marginal distributions. Let $T_{\Delta}(\sigma,\sigma^{\prime})=\text{Sign}(\Delta)_{\sigma,\sigma^{\prime}}$ be the agreement function where $\text{Sign}(x)=1$ if $x>0$ and 0 otherwise. Intuitively, $T_{\Delta}(\sigma,\sigma^{\prime})=1$ if the pair of signals $(\sigma,\sigma^{\prime})$ appears more often on the same question than on distinct questions.
Then, the mechanism repeatedly draws a bonus question $j$ and a penalty question $k$. For an individual score $S^{CA}_{i}$, Alice gets 1 if both agents “agree” on $j$ and gets $-1$ if Alice’s response on $j$ “agrees” with Bob’s response on $k$, where the agreement is determined by $T_{\Delta}$, i.e. $S^{CA}_{i}=T_{\Delta}(\hat{X}_{a,j},\hat{X}_{b,j})-T_{\Delta}(\hat{X}_{a,j},\hat{X}_{b,k})$. Each bonus question can be paired with $n-1$ possible penalty questions. Therefore, the final score of CA is the average of $n\cdot(n-1)$ individual scores.
Intuitively, the CA mechanism works by encouraging correlations on the same question and penalizing correlations on distinct questions. Consequently, if Alice always reports the majority signal, the frequency of agreements on the bonus question will be identical to the frequency of agreements on the penalty questions, which results in a score of 0.
Previous results suggest that if the information structure $\Pr(X_{a},X_{b})$ is known (or accurately estimated), CA is truthful (Shnayder et al., 2016). [^3] In fact, under these conditions, CA is informed-truthful.
Is CA SD-truthful?  The answer depends on how we implement the mechanism. In general, the original implementation of CA described above is not SD-truthful. However, we first present an intermediate result focusing on scoring Alice with an individual CA score $S^{CA}_{i}$.
Proposition 4.3.
However, as we will soon discuss in Section 5, scoring agents via a binary score is not sensitive. In practice, we need to take the average over these samples as the final score. Our second observation warns that we should be careful about this averaging process: repeatedly using the same question as the bonus or the penalty question will sabotage the SD-truthfulness of the mechanism.
@@
Suppose $n=2$, meaning that Alice knows each of the two questions she answers will be used once as the bonus question and once as the penalty question. Her final score is the average of two samples, where each sample takes values $S^{CA}_{i}\in\{-1,0,1\}$. Consequently, the final score can take five possible values: $S^{CA}\in\{-1,-0.5,0,0.5,1\}$. For SD-truthfulness, a necessary condition is that for any strategy $\theta$, we must have $\Pr(S^{CA}(\tau)=-1)\leq\Pr(S^{CA}(\theta)=-1)$. However, we provide a counterexample in the binary-signal setting showing that this condition does not hold.
Let $J_{\sigma,\sigma^{\prime}}=\Pr(X_{a}=\sigma,X_{b}=\sigma^{\prime})$ denote the joint distribution, and let $M^{a}_{\sigma}=\Pr(X_{a}=\sigma)$ and $M^{b}_{\sigma}=\Pr(X_{b}=\sigma)$ denote the marginal distribution of Alice and Bob’s signals, respectively, where $\sigma,\sigma^{\prime}\in\{0,1\}$. Suppose the signals are positively correlated, i.e. $T_{\Delta}$ is a diagonal matrix. Alice receives a score of $-1$ if and only if 1) she reports different signals on the two questions, and 2) the agents disagree on both questions. Under truth-telling, this occurs with a positive probability: $\Pr(S^{CA}(\tau)=-1)=2J_{0,1}J_{1,0}$. However, if Alice always reports the same signal, she can avoid this outcome entirely. This suggests that an uninformative strategy $\mu(\sigma)=0$ for any $\sigma$ is not stochastically dominated by truth-telling.
Consequently, while the original implementation of the CA mechanism achieves the highest sensitivity, it is not SD-truthful. We will discuss how to implement CA (and other mechanisms) to achieve SD-truthfulness without destroying its sensitivity.
@@ 5 A Rounding Reduction For Stochastically Dominant-Truthfulness
In this section, we introduce a straightforward rounding reduction that can transform a truthful mechanism into an SD-truthful one. However, this approach often comes at a huge cost of sensitivity. To address this, we propose a more refined application of the rounding technique that partially restores the sensitivity of the original mechanism.
@@ 5.1 Direct Rounding
The direct rounding reduction of a truthful mechanism $\mathcal{M}$ with bounded scores is SD-truthful.
The proof is straightforward. First, note that the rounded score $\tilde{S}^{\mathcal{M}}$ is a Bernoulli random variable with success probability $\mathbb{E}[\lambda^{\mathcal{M}}]$. For any truthful mechanism $\mathcal{M}$, truth-telling maximizes the expected score, which in turn maximizes the expected success probability of $\tilde{S}^{\mathcal{M}}$. Since $\tilde{S}^{\mathcal{M}}$ takes only binary values, maximizing the probability of receiving a score of 1 ensures that truth-telling stochastically dominates the score distribution of any other strategy. Therefore, scoring agents according to $\tilde{S}^{\mathcal{M}}$ is SD-truthful.
@@
Proof of 5.2.
[…]
Intuitively, this ratio is usually large. Consider the output agreement mechanism, whose final score is the average of $n$ i.i.d. binary scores, i.e. a binomial distribution with a success probability of exactly $m(e)$. A standard result for the binomial variable suggests that the standard deviation of OA is $\sqrt{\frac{m(e)(1-m(e))}{n}}$. Therefore, the ratio between the sensitivity of OA and its direct rounding reduction is $\sqrt{n}$. This great difference highlights a significant drawback of direct rounding: while it ensures stochastic dominance, it significantly reduces the sensitivity of the mechanism by preventing the score from averaging out over multiple questions.
@@ 5.2 Partition Rounding
To reduce the variance of the final score and improve sensitivity, we need to make better use of the information from the $n$ questions. In particular, if the final score of the peer prediction mechanism $S^{\mathcal{M}}$ is the average of $K$ individual scores $S^{\mathcal{M}}_{i}$, we can apply direct rounding to each individual score and then average the results after rounding. The following lemma proves the feasibility of this idea.
@@
We defer the proof, which follows from a straightforward coupling argument, to Section A.2. Lemma 5.3 motivates the idea of partition rounding, where the set of $n$ questions is divided into $K$ disjoint subsets. Each subset is then used to compute an individual score, which is subsequently rounded using the direct rounding reduction. The partition process guarantees that each individual score is i.i.d. because the questions are assumed to be i.i.d. and Alice’s strategy is task-independent. This design prevents the failure observed in the original CA mechanism, which suggests that each question should be used only once when computing an individual score.
@@
The partition rounding reduction of a truthful mechanism $\mathcal{M}$ with bounded scores is SD-truthful.
The proposition follows directly from 5.1 and Lemma 5.3.
We now show that the partition rounding reduction can partially recover the sensitivity of the original mechanism. The key observation is that the final score under partition rounding is the average of a binomial variable, i.e. $\hat{S}^{\mathcal{M}}(e)=\frac{1}{K}\mathrm{Bin}(K,\lambda^{\mathcal{M}}(e))$, where $\lambda^{\mathcal{M}}(e)=\frac{m^{\mathcal{M}}-S^{\mathcal{M}}_{\inf}}{S^{\mathcal{M}}_{\sup}-S^{\mathcal{M}}_{\inf}}$ with $m^{\mathcal{M}}$ being the expectation of the individual score. Putting this into Definition 3 gives us the sensitivity of the partition rounding reduction

### v2

@@ 3.2 Sensitivity
In practice, an effective scoring mechanism should meaningfully evaluate the information contained in agents’ reports and reward more informative reports with higher scores. This motivates an additional design criterion, orthogonal to truthfulness, known as sensitivity (Zhang and Schoenebeck, 2023a).
To motivate the concept of sensitivity, we extend the model to incorporate agents’ effort levels. This extension captures how the mechanism responds to marginal changes in information at equilibrium, i.e. how the score varies when all other agents exert full effort while one agent adjusts her effort. In particular, on each question, we assume Alice exerts effort with probability $e\in[0,1]$, while Bob always exerts effort and reports truthfully. When Alice exerts effort, the signal pair $(X_{a},X_{b})$ is drawn from the joint distribution $\Pr(X_{a},X_{b})$; in the extreme case where Alice exerts no effort, the signals are independently drawn from the marginals $\Pr(X_{a})\Pr(X_{b})$. That is, exerting effort produces a signal $X_{a}$ that is correlated with $X_{b}$ according to the joint distribution, while shirking yields an uninformative guess drawn from the prior $\Pr(X_{a})$. After observing the signal, Alice will report truthfully. Then, Alice’s score under mechanism $\mathcal{M}$ becomes a function of $e$, denoted as $S^{\mathcal{M}}(e)$.
Definition 3.3.
@@
Sensitivity measures how well a mechanism’s expected score responds to changes in effort—and thus, changes in information. Consequently, a trivial mechanism that always assigns a score of 1, though truthful, has zero sensitivity, as its output is unaffected by effort.
Prior work (Zhang and Schoenebeck, 2023a) shows that when the score is normally distributed and agents are rewarded based on the ranking of scores, sensitivity is a sufficient statistic for budgetary efficiency: mechanisms with higher sensitivity can elicit the same equilibrium effort level at a lower budgetary cost. Moreover, Xu et al. (2024) establish a one-to-one correspondence between sensitivity and measurement integrity (Burrell and Schoenebeck, 2023), implying that mechanisms with higher sensitivity generate scores more closely aligned with the true quality of responses. Together, these findings highlight sensitivity as a practically meaningful and theoretically grounded property of peer prediction mechanisms, motivating our choice to use it as an additional dimension for comparison.
@@ 4 Prior Mechanisms
To the best of our knowledge, no existing peer prediction mechanism naturally achieves SD-truthfulness without imposing strong assumptions on the information structure. In this section, we introduce three exemplary multi-task peer prediction mechanisms and explain why they are not naturally SD-truthful to motivate our methods.
@@ 4.1 Output Agreement (OA)
The output agreement (OA) mechanism scores Alice based on the fraction of questions on which both agents agree. The final score of OA is the average of $n$ independent binary individual scores, i.e. $S^{OA}=\frac{1}{n}\sum_{i\in[n]}S_{i}^{OA}$ where $S_{i}^{OA}=1$ if both agents’ reports agree on question $i$ and 0 otherwise. However, OA encourages over-reporting of the majority signal, which is truthful only under a strong assumption.
Definition 4.1.
@@
OA is truthful if agents’ signals are self-dominating.
In fact, OA is also SD-truthful when signals are self-dominating. This is because each individual score $S^{OA}_{i}$ takes only two values, meaning that $S^{OA}_{i}(\tau)$ FOSD $S^{OA}_{i}(\theta)$ if and only if $\Pr(S^{OA}_{i}(\tau)=1)\geq\Pr(S^{OA}_{i}(\theta)=1)$. Furthermore, the average of multiple i.i.d. scores that are each SD-truthful remains SD-truthful (see Section 5.2). Consequently, for OA, SD-truthfulness reduces to truthfulness.
@@ 4.2 Peer Truth Serum (PTS)
Self-dominance is usually considered a strong assumption. (Radanovic et al., 2016) propose the Peer Truth Serum (PTS) which incentivizes truth-telling under a weaker assumption. [^2] In the binary signal setting, self-prediction is equivalent to positive correlation, which is a strictly weaker condition than self-dominance.
Definition 4.3.
@@
PTS assumes a symmetric, publicly known prior $R(\sigma)=\Pr(X_{a}=\sigma)=\Pr(X_{b}=\sigma)$. It repeatedly samples questions without replacement, and assigns an individual score $S^{PTS}_{i}=1/R(\sigma)$ to Alice if $\hat{X}_{a,i}=\hat{X}_{b,i}=\sigma$, and 0 otherwise. The final score of PTS is the average of all $S^{PTS}_{i}$.
@@
PTS is truthful if agents’ signals are self-predicting.
Is PTS SD-truthful?  The short answer is no. Each individual score $S^{PTS}_{i}$ takes at most $|\Sigma|+1$ values, with the maximum being $S_{\text{max}}=1/R(\sigma_{\min})$, where $\sigma_{\min}$ is the least likely signal under the prior. A necessary condition for PTS to be SD-truthful is that truth-telling maximizes $\Pr\left(S^{PTS}_{i}(\tau)=S_{max}\right)$, which is the probability of Alice agreeing with Bob on signal $\sigma_{min}$. While truth-telling, this probability is the joint distribution $\Pr(X_{a}=\sigma_{min},X_{b}=\sigma_{min})$. However, if Alice uses an uninformative strategy $\mu$ that always reports $\sigma_{\min}$, this probability increases to $\Pr(X_{b}=\sigma_{\min})$. Therefore, although $\mu$ decreases the expected score, its distribution is not stochastically dominated by truth-telling, meaning that there exists an increasing utility function under which Alice strictly prefers $\mu$ to $\tau$.
@@ 4.3 Correlated Agreement (CA)
The correlated agreement (CA) mechanism (Shnayder et al., 2016) takes two questions to compute an individual score which has three possible values: -1, 0, and 1. Suppose the joint distribution $\Pr(X_{a},X_{b})$ is known. The CA mechanism first computes the delta matrix $\Delta_{\sigma,\sigma^{\prime}}=\Pr(X_{a}=\sigma,X_{b}=\sigma^{\prime})-\Pr(X_{a}=\sigma)\Pr(X_{b}=\sigma^{\prime})$, which is the difference between the joint distribution and the product of marginal distributions. Let $T_{\Delta}(\sigma,\sigma^{\prime})=\text{Sign}(\Delta)_{\sigma,\sigma^{\prime}}$ be the agreement function where $\text{Sign}(x)=1$ if $x>0$ and 0 otherwise. Intuitively, $T_{\Delta}(\sigma,\sigma^{\prime})=1$ if the pair of signals $(\sigma,\sigma^{\prime})$ appears more often on the same question than on distinct questions.
Then, CA repeatedly samples a bonus question $j$ and a penalty question $k$. For an individual score $S^{CA}_{i}$, Alice gets 1 if both agents agree on $j$ and gets 1-1 if her response on $j$ agrees with Bob’s response on $k$, i.e. $S^{CA}_{i}=T_{\Delta}(\hat{X}_{a,j},\hat{X}_{b,j})-T_{\Delta}(\hat{X}_{a,j},\hat{X}_{b,k})$. Since each bonus question can be paired with n1n-1 penalties, the final score is the average over n(n1)n(n-1) individual scores.
Intuitively, the CA mechanism works by encouraging correlations on the same question and penalizing correlations on distinct questions. Consequently, if Alice always reports the majority signal, the frequency of agreements on the bonus question will be identical to the frequency of agreements on the penalty questions, which results in a score of 0. Prior work shows that when $\Pr(X_{a},X_{b})$ is known (or accurately estimated), CA is truthful—and in fact, informed truthful (Shnayder et al., 2016).
Is CA SD-truthful?  The answer depends on the implementation. The following example illustrates that the original implementation of CA described above is not SD-truthful.
@@
Suppose $n=2$, so that Alice knows each of the two questions she answers will be used once as the bonus question and once as the penalty question. Her final score is the average of two individual scores, each with $S^{CA}_{i}\in\{-1,0,1\}$. Thus, the final score has five possible values: $S^{CA}\in\{-1,-0.5,0,0.5,1\}$. A necessary condition for SD-truthfulness is that $\Pr(S^{CA}(\tau)=-1)\leq\Pr(S^{CA}(\theta)=-1)$ for any strategy $\theta$. However, we show that this does not hold in general.
Suppose the signals are positively correlated, i.e. $T_{\Delta}$ is a diagonal matrix. Alice receives a score of 1-1 if and only if she reports different signals on the two questions and Bob disagrees on both. Under truth-telling, this happens with probability $\Pr(S^{CA}(\tau)=-1)=2\Pr(X_{a}=0,X_{b}=1)\Pr(X_{a}=1,X_{b}=0)$. However, if Alice always reports the same signal (e.g., $\mu(\sigma)=0$ for any $\sigma$), she can avoid this outcome entirely. Therefore, $\mu$ is not stochastically dominated by $\tau$.
This warns that repeatedly using the same question as the bonus or the penalty question will sabotage the SD-truthfulness of CA.
@@ 5 A Rounding Reduction For Stochastically Dominant-Truthfulness
We introduce a straightforward rounding reduction that maps any truthful mechanism to an SD-truthful one. However, this approach often comes at a huge cost of sensitivity. To address this, we propose a more refined rounding method that partially restores the original mechanism’s sensitivity.
@@ 5.1 Direct Rounding
The direct rounding reduction of a truthful mechanism with bounded scores is SD-truthful.
Intuitively, because direct rounding always outputs a Bernoulli score, SD-truthfulness simply implies that truth-telling should maximize the expected success probability. But the expected success probability is just the success probability. By our design of $\lambda^{\mathcal{M}}$, the success probability is a linear function of the final score $S^{\mathcal{M}}$, which proves equivalence between SD-truthfulness and truthfulness.
@@
This ratio is typically large. For example, the final score of OA is an averaged Binomial $\mathrm{Bin}(n,m^{OA}(e))$, whose standard deviation is $\sqrt{m^{OA}(e)(1-m^{OA}(e))/n}$. Thus, the sensitivity ratio between OA and its direct rounding is $\sqrt{n}$. This illustrates a key drawback of direct rounding: it greatly reduces the sensitivity by eliminating the benefit of averaging over multiple questions.
@@ 5.2 Partition Rounding
To reduce variance of and improve sensitivity, we aim to better leverage the information from all $n$ questions. In particular, if the final score of $S^{\mathcal{M}}$ is the average of $K$ independent individual scores $S^{\mathcal{M}}_{i}$, we can apply direct rounding to each $S^{\mathcal{M}}_{i}$ and take the average. This motivates the idea of partition rounding. We first partition the $n$ questions into $K$ disjoint subsets, using each to compute an individual score. Each score is then directly rounded as in Section 5.1, and the final score $\hat{S}^{\mathcal{M}}$ is the average of the $K$ rounded scores. The following lemma proves the feasibility of this idea.
@@
The partition rounding reduction of a truthful mechanism with bounded scores is SD-truthful.
We defer the proof of Lemma 5.3, which follows from a straightforward coupling argument, to Section F.3. 5.4 then follows directly from this lemma and the SD-truthfulness of direct rounding.
We now show that the partition rounding reduction can partially recover the sensitivity of the original mechanism. The key observation is that the final score under partition rounding is the average of a binomial variable, i.e. $\hat{S}^{\mathcal{M}}(e)=\frac{1}{K}\mathrm{Bin}(K,\lambda^{\mathcal{M}}(e))$, where $\lambda^{\mathcal{M}}(e)=\frac{m^{\mathcal{M}}-S^{\mathcal{M}}_{\inf}}{S^{\mathcal{M}}_{\sup}-S^{\mathcal{M}}_{\inf}}$ with $m^{\mathcal{M}}$ being the expectation of the individual score. Putting this into Definition 3.3 gives us the sensitivity of the partition rounding reduction